What is the pressure exerted by 8.47 g of krypton with a volume of 1477.2 mL at a temperature of -20.2oC?
1
Expert's answer
2019-10-11T09:39:29-0400

Use ideal gas Law


"PV = nRT"

As n=m/M, then



"PV =\frac{m}{M} RT"

"P\times1.4772 L =\frac{8.47 g}{83.80\frac{g}{mol}}\times0.082\frac{L\times atm}{K\times mol}\times (-20.2+273.15) K"

"P=1.419 atm"